# Reinforcement Learning with Hidden Markov Models for Discovering Decision-Making Dynamics

### v1

@@ 2.4 Parameter inference and model evaluation
In addition to point estimation, assessing the variability of parameter estimates is essential for statistical inference. We use nonparametric bootstrapping inference, where the reward learning process for each subject $(\mathbf{S}_{i,[1:T]},\mathbf{A}_{i,[1:T]},\mathbf{R}_{i,[1:T]})$ is considered as a unified entity and resampled with replacement. We construct $95\%$ bootstrap confidence intervals under the normal distribution (i.e., $\mbox{MLE}\pm 1.96\times\mbox{bootstrap se.}$.) in simulation studies and data application. Since the distribution of $\pi_{1}$ is left-skewed when the true $\pi_{1}>0.5$ and right-skewed when the true $\pi_{1}<0.5$, we apply a logit transformation to $\pi_{1}$ before using a normal approximation to construct confidence intervals.
One way to evaluate the performance of a model is by utilizing the $K$-fold cross-validation and computing the out-of-sample observed joint log-likelihood. Recall that $a_{ijt}(\bm{\theta})=P\left(U_{it}=j,\mathbf{A}_{i,[1:t]};\bm{\theta}\right)$, for subject $i$, the observed likelihood is
$P(\mathbf{A}_{i,[1:T]};\bm{\theta})=\sum_{j=0}^{1}P\left(U_{iT}=j,\mathbf{A}_{i,[1:T]};\bm{\theta}\right)=\sum_{j=0}^{1}a_{ijT}(\bm{\theta}).$
Therefore, the $K$-fold cross-validation score is
@@ 3 Simulation studies
We conducted extensive simulation studies to assess the finite-sample performance of the proposed method. For the RL model, we considered a continuous state space $\mathcal{S}=[0,1]$ and a two-arm action space $\mathcal{A}=\{0,1\}$. Denote the distribution for generating reward $R_{it}$ conditional on $A_{it}$ and $S_{it}$ as $f_{r}(R_{it}\mid A_{it},S_{it})$. We considered the reward-generating distribution as $f_{r}(R_{it}\mid A_{it}=0,S_{it}=s)\sim\mbox{Bernoulli}(p_{0}(1-s))$, $f_{r}(R_{it}\mid A_{it}=1,S_{it}=s)\sim\mbox{Bernoulli}(p_{1}s)$. We let $p_{0}=0.5$, $p_{1}=1$. The rewards at different time points were assigned independently. The states were independently generated by uniform distribution $\mbox{U}(0,1)$. We set learning rate $\beta=0.05$, reward sensitivity $\rho=4$, and intercept $\nu_{1}=0$. We choose a linear basis $\phi^{\top}(a,s)=(I(a=0),I(a=1))\otimes(1,s)$ and assume that the initial coefficient has the form $\bm{\alpha}=\alpha\cdot(1,-1,0,1)$, where we set $\alpha=2$. We considered two HMM cases: switching between engaged and lapse (Case I) and always engaged (Case II). For Case I, suppose $\pi_{1}=0.8$, the transition probabilities are piecewise constant functions with $\zeta_{0t}=-1.5\cdot(1+I(T/2\leq t<T))$ and $\zeta_{1t}=2\cdot(1+I(T/2\leq t<T))$. For Case II, there is no learning strategy switching.
@@
We generated 200 replicates with sample sizes $n\in\{100,200\}$ and the number of trials $T\in\{100,200\}$. We estimated the parameters of interest, $\bm{\theta}=\left\{\beta,\rho,\nu_{1},\alpha,\pi_{1},\bm{\zeta}_{0},\bm{\zeta}_{1}\right\}$, using the proposed EM algorithm with a fused Lasso penalty. The regularization parameters $(\lambda_{0},\lambda_{1})$ were selected by maximizing the $5$-fold cross-validation scores in (13), where the values of the regularization parameters were chosen from +$\infty$+ ∞ to 00, at which the solution path of generalized Lasso changes slope (Arnold and Tibshirani, 2016).
For Case I, we used $50$ bootstrap samples to construct $95\%$ bootstrap confidence intervals for the parameters under the normal distribution. The evaluation of $\bm{\zeta}_{0}$ and $\bm{\zeta}_{1}$ was conducted at $t=T/4$ and $t=3T/4$, respectively. The simulation results of the parameter estimate from $200$ replicates were given in Table 1. The estimate bias (Bias), standard deviation (SD), average bootstrap standard error of the $50$ bootstrap samples (SE), and coverage probability of the $95\%$ bootstrap confidence intervals (CP) were reported. Furthermore, we compared three approaches: (i) our proposed method (RL-HMM), (ii) the RL-HMM model with fixed transition probabilities over trials (RL-HMM-fixed), and (iii) the RL only model without learning strategy switching (RL-only).
@@
Table 1 demonstrates that when subjects’ decision-making involves a mixture of two learning strategies, the RL-only model tends to underestimate the reward sensitivity ($\rho$) and the initial coefficient of Q values ($\alpha$) in comparison to the RL-HMM. Additionally, when the transition probabilities are time-varying, employing a time-invariant RL-HMM results in a larger absolute bias for most of the parameters. Meanwhile, for our proposed RL-HMM, the bias of most parameters decrease when $n$ and $T$ increase. Furthermore, the estimates of $\zeta_{0t}$ and $\zeta_{1t}$ are biased due to the presence of generalized lasso penalty terms, indirectly causing bias in $\rho$. The coverage probabilities of all parameter estimates are close to the nominal level ($95\%$) for all three cases.
@@
Finally, in both Case I and Case II, we presented the 5-fold cross-validation (CV) scores for the 200 replicates using boxplots in Figure 1 (a) and (b). The oracle (i.e., parameters are known) is compared with the RL-HMM, RL-HMM-fixed, and the RL model. In Figure 1 (a) for Case I, we observed that the RL-HMM demonstrates nearly the same performance as the oracle and outperforms RL-HMM-fixed, with RL showing the least favorable CV performance. In Figure 1 (b) for Case II, all models exhibit similar CV performance. On the other hand, recalling the estimated posterior probability of subject $i$ being in the ‘engaged’ learning strategy at trial $t$ as $\widehat{\gamma}_{i1t}=P\left(U_{it}=1\mid\mathbf{A}_{i},\widehat{\bm{\theta}}\right)$, we estimate the learning strategy at trial $t$ for subject $i$ by $\widehat{U}_{it}=I(\widehat{\gamma}_{i1t}\geq 0.5)$. We present the averaged accuracy for estimating $U_{it}$ (i.e., $(nT)^{-1}\sum_{i=1}^{n}\sum_{t=1}^{T}I(\widehat{U}_{it}=U_{it})$) for the 200 replicates using boxplots in Figure 1 (c) and (d). In Figure 1 (c) for Case I, the RL-HMM outperforms RL-HMM-fixed, with RL showing the lowest accuracy because the RL model classifies every $U_{it}$ as ‘engaged’. In Figure 1 (d) for Case II, RL-HMM estimates more than $99.9\%$ of $U_{it}$ correctly for most of the replicates, demonstrating the robustness of RL-HMM when the model is misspecified.
@@ 4.1 Probabilistic reward task and EMBARC study
In the EMBARC study (Trivedi et al., 2016), the PRT, as introduced by Pizzagalli et al. (2005), is a computer-based experiment used to assess subjects’ ability to adapt their behavior in response to rewards. During each trial of the task, participants are presented with a cartoon face displaying one of the two stimuli (i.e., a face with a short mouth or with a long mouth). The objective for participants is to identify the type of stimulus in each trial. Crucially, the PRT provides a higher frequency of rewarding participants for correctly identifying faces with a short mouth, designating it as the ‘rich reward’, while providing less frequent rewards for correctly identifying faces with a long mouth, referred to as the ‘lean reward’. Participants receive verbal instructions outlining the task’s goal, which is to maximize the total rewards. It is essential that participants understand that not all correct responses will be rewarded. To optimize their rewards, participants should focus on providing the correct response, regardless of the face associated with the higher reward. However, the difference in mouth length between the short and long mouths is intentionally kept small. The subtle difference in mouth length often leads to difficulties in accurately perceiving the stimuli presented, resulting in a tendency to prioritize states with higher rewards over those that are more accurate.
In the EMBARC study, each PRT session comprised 200 trials, divided into two blocks of 100 trials, with a 30-second break separating the blocks. In each block, 40 correct trials were followed by reward feedback. Correct identification of the rich stimulus was associated with significantly more positive feedback (30 out of 40 trials) compared to correct identification of the lean stimulus (10 out of 40 trials). The EMBARC study was designed to discern differences in reward learning abilities between patients diagnosed with Major Depressive Disorder (MDD) and a healthy control group. The PRT experiments involved 40 subjects from the control group and 168 subjects from the MDD group. Each subject in the MDD group had two PRT sessions, one session at the baseline before treatment (week 0) and one session after one week of treatment (week 1). Subjects in the control group also took two sessions in week 0 and week 1 (no treatment at both times). More details about the EMBARC study and the PRT experiment can be found in Guo et al. (2023).
@@ 4.2 Model fitting and results
We now apply the proposed methodology to our motivating PRT data. Let $\mathcal{S}=\{0,1\}$, where 00 and $1$ represent the type of stimulus (rich versus lean); $\mathcal{A}=\{0,1\}$, where 00 and $1$ represent the subject choosing lean and rich stimulus, respectively. We choose the basis $\phi(a,s)=(\delta_{00},\delta_{10},\delta_{01},\delta_{11})^{\top}$, where $\delta_{uv}=I(a=u,s=v)$ $u,v\in\{0,1\}$. We assume no initial preference for one stimulus state over the other at the beginning of each session. To ensure this, we specify the initial coefficients as $\bm{\alpha}=(\alpha,0,0,\alpha)$ and set the RL intercept values as $\nu_{0}=\nu_{1}=0$. In a preliminary analysis, we observed that the learning patterns may change between the first and second block. To mitigate any potential bias, we analyze the PRT data from the first block in this paper. We started by fitting a separate RL model for each session, using the method outlined in Section 2.1, and excluded sessions with a learning rate less than $10^{-3}$. We fitted our models for the MDD group that contains subjects diagnosed with MDD at week 0 (pre-treatment) and the control group with two repeated measurements, respectively. Based on the $5$-fold cross-validation likelihood score, we choose an RL-HMM with fixed transition probabilities over trials. This model has a higher score than an RL-HMM with time-varying transition probabilities or an RL-only model without learning strategy switching. The values of $C_{01}$ are $2.8\%$ for the MDD group and $1.3\%$ for the control group. Similarly, the values of $C_{11}$ are $99.2\%$ for the MDD group and $99.7\%$ for the control group. For inference, a nonparametric bootstrap was applied to generate 200 resampling sets for each group. Bootstrap standard errors and $95\%$ bootstrap confidence intervals are shown in Table 3.
@@
Table 3 reveals that the differences in the learning rate ($\beta$), reward sensitivity ($\rho$), and the initial probability of being in the ‘engaged’ learning strategy ($\pi_{1}$) between the MDD group and the control group are not statistically significant. This finding contrasts with previous research, such as Huys et al. (2013) and Guo et al. (2023), in which they reported a significantly lower reward sensitivity in the MDD group compared to the healthy control group in RL-only models. On the other hand, the MDD group presents a significantly lower initial coefficient ($\alpha$) compared to the control group. This indicates that the MDD group has a lower probability of selecting the correct state (i.e., distinguishing between long and short mouths) at the beginning of the task.
@@
Recall that the posterior probability of the subject $i$ being in the ‘engaged’ learning strategy at trial $t$ is denoted as ${\gamma}_{i1}(t):={\gamma}_{i1t}=P\left(U_{it}=1\mid\mathbf{A}_{i},\bm{\theta}\right)$. We refer to this quantity as the individual engagement probability at trial $t$. Figure 2(a) presents the maximum likelihood estimates of ${\gamma}_{i1}(t)$ for four randomly selected MDD patients which shows a general decreasing tendency of being engaged but with heterogeneous patterns between patients. We define the group engagement rate at trial $t$ as the average of the individual engagement probabilities in the group (i.e. $n^{-1}\sum_{i=1}^{n}{\gamma}_{i1}(t)$). The maximum likelihood estimates and the $95\%$ pointwise bootstrap confidence intervals for the MDD and control groups engagement rates are shown in Figure 2(b). We observe that the engagement rates for MDD and control groups are similar in the first 3/4 of the task. However, the MDD group engagement rate decreases and becomes smaller in the last 1/4 of the task compared to the control group engagement rate. This implies that individuals in the MDD group are more likely to be in the ‘lapse’ state after many trials in contrast to the control group. This finding suggests an alternative explanation for the reward processing abnormalities in MDD; it may be attributed to a reduced duration of making decisions using RL rather than a diminished sensitivity to rewards.
@@
We define the individual engagement score as $\mbox{ES}_{i}(\mathcal{T})=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\log(\gamma_{i1t}/(1-\gamma_{i1t}))$. This score quantifies the average (logit transformation of) individual engagement probabilities over trials in the trial set $\mathcal{T}$. To demonstrate the utilities of engagement scores, we consider its association with a commonly used clinical measure of depression severity, the Hamilton Depression Rating Scale (HDRS) (Hamilton, 1960). In EMBARC, the HDRS was administered by trained clinicians to assess depression symptoms in different domains at the same time when the PRT was collected. One of items in HDRS assesses concentration (referred to as distraction level) on a liker scale of zero to three, where a higher value indicates greater difficulty in concentrating. In Figure 3(a), we present boxplots of individual engagement scores at different distraction levels (ranging from 0 to 3) in four periods of the task for patients in the MDD group. We observe that as the distraction level increases, the individual engagement score decreases. This effect is more pronounced in the first 3/4 of the trials. Furthermore, we performed a regression analysis of distraction levels on the individual engagement score using generalized linear models with binomial link functions. We found that the $p$-values for the coefficients are $0.004$, $0.043$, $0.013$, and $0.090$ for the four periods of the tasks, respectively. This implies that a lack of engagement, particularly at the beginning of the task, is associated with difficulties in concentration.
@@
Define $\widehat{U}_{it}=I(\widehat{\gamma}_{i1t}\geq 0.5)$ to be the predicted type of learning strategy for the subject $i$ at trial $t$. The correct response rates (i.e. the accurate identification of short and long mouth stimuli) are $80.9\%$ and $48.4\%$ for the ‘engaged’ and ‘lapse’ state within the MDD group, and $84.3\%$ and $55.3\%$ for the control group. This suggests that the MDD group has a lower correct response rate in the ‘engaged’ state, and the correct response rate in the ‘lapse’ state is close to random guessing in both groups. Additionally, Iigaya et al. (2018) showed that different learning strategies may result in different lengths of intertrial intervals (ITIs), which is the response time between the presence of a stimulus and the decision making in a trial. Furthermore, we obtain the smoothed ITIs over trials for the two learning strategies for two groups separately using local polynomial regression (loess function in R). The mean estimates and their $95\%$ confidence intervals for the mean curves are displayed in Figure 3(b). It reveals that, on average, the ‘engaged’ learning strategy requires more time for individuals to make decisions compared to the ‘lapse’ learning strategy. Meanwhile, the control group, on average, takes less time to make decisions than the MDD group.

### v2

@@ 2.4 Parameter inference and model evaluation
In addition to point estimation, assessing the variability of parameter estimates is essential for statistical inference. We employ nonparametric bootstrapping inference, treating the reward learning process for each subject $\mathcal{H}_{i,T+1}=(\mathbf{S}_{i},\mathbf{A}_{i},\mathbf{R}_{i})$ as a unified entity, which is then resampled with replacement. Hence, the bootstrap method is deemed valid as $\mathcal{H}_{i,T+1}$ $(i=1,\dots,n)$ are assumed to be independent and identically distributed (i.i.d.) when $n$ is large. We construct $95\%$ bootstrap confidence intervals under the normal distribution (i.e., $\mbox{MLE}\pm 1.96\times\mbox{bootstrap se.}$.) in simulation studies and data application. Since the distribution of $\pi_{1}$ is left-skewed when the true $\pi_{1}>0.5$ and right-skewed when the true $\pi_{1}<0.5$, we apply a logit transformation to $\pi_{1}$ before using a normal approximation to construct confidence intervals.
One way to evaluate the performance of a model is by utilizing the $K$-fold cross-validation and computing the out-of-sample observed joint log-likelihood. For subject $i$, recall the observed likelihood takes form: $P(\mathbf{A}_{i,[1:T]};\bm{\theta})=\sum_{j=0}^{1}a_{ijT}(\bm{\theta})$. Thus, the $K$-fold cross-validation score is
@@ 3 Simulation studies
We conducted extensive simulation studies to assess the finite-sample performance of the proposed method. For the RL model, we considered a continuous state space $\mathcal{S}=[0,1]$ and a two-arm action space $\mathcal{A}=\{0,1\}$. We considered the reward-generating distribution as $p(R_{it}\mid A_{it}=a,S_{it}=s)\sim\mbox{Bernoulli}\big{(}p_{a}\{as+(1-a)(1-s)\}\big{)}$. We let $p_{0}=0.5$, $p_{1}=1$. The rewards at different time points were assigned independently. The states were independently generated by uniform distribution $\mbox{U}(0,1)$. We set learning rate $\beta=0.05$, reward sensitivity $\rho=4$, and intercept $\nu_{1}=0$. We choose a linear basis $\phi^{\top}(a,s)=(I(a=0),I(a=1))\otimes(1,s)$ and assume that the initial coefficient has the form $\bm{\alpha}=(\alpha/\rho)\cdot(1,-1,0,1)$, where we set $\alpha=2$. We considered two HMM cases: switching between engaged and lapse (Case I) and always engaged (Case II). For Case I, suppose $\pi_{1}=0.8$, the transition probabilities are piecewise constant functions with $\zeta_{0t}=-1.5\cdot(1+I(T/2\leq t<T))$ and $\zeta_{1t}=2\cdot(1+I(T/2\leq t<T))$. For Case II, there is no decision-making strategy switching.
@@
We generated 200 replicates with sample sizes $n\in\{100,200\}$ and the number of trials $T\in\{100,200\}$. We estimated the parameters of interest, $\bm{\theta}=\left\{\beta,\rho,\nu_{1},\alpha,\pi_{1},\bm{\zeta}_{0},\bm{\zeta}_{1}\right\}$, using the proposed EM algorithm with a fused Lasso penalty. The regularization parameters $(\lambda_{0},\lambda_{1})$ were selected by maximizing the $5$-fold cross-validation scores in (10), where the values of the regularization parameters were chosen from +$\infty$+ ∞ to 00, at which the solution path of generalized Lasso changes slope (Arnold and Tibshirani,, 2016).
@@
For Case I, we used $50$ bootstrap samples to construct $95\%$ bootstrap confidence intervals for the parameters under the normal distribution. The evaluation of $\bm{\zeta}_{0}$ and $\bm{\zeta}_{1}$ was conducted at $t=T/4$ and $t=3T/4$, respectively. The simulation results of the parameter estimate from $200$ replicates were given in Table 1. The estimate bias (Bias), standard deviation (SD), average bootstrap standard error of the $50$ bootstrap samples (SE), and coverage probability of the $95\%$ bootstrap confidence intervals (CP) were reported. Furthermore, we compared three approaches: (i) our proposed method (RL-HMM), (ii) the RL-HMM model with fixed transition probabilities over trials (RL-HMM-fixed), and (iii) the RL model without decision-making strategy switching (RL-only). Table 1 demonstrates that when subjects’ decision-making involves a mixture of two decision-making strategies, the RL-only model tends to underestimate the reward sensitivity ($\rho$) and the initial coefficient of Q values ($\alpha$) in comparison to the RL-HMM. Additionally, when the transition probabilities are time-varying, employing RL-HMM-fixed results in a larger absolute bias for most of the parameters. Meanwhile, for our proposed RL-HMM, the bias of most parameters decrease when $n$ and $T$ increase. Furthermore, the estimates of $\zeta_{0t}$ and $\zeta_{1t}$ are biased due to the presence of generalized lasso penalty terms, indirectly causing bias in $\rho$. The coverage probabilities of all parameter estimates are close to the nominal level ($95\%$) for all three cases. In the Supplementary Materials Section S.2, we show that 200 replicates and 50 bootstrap samples are sufficient to provide accurate point estimates and confidence intervals.
@@
Finally, in both Case I and Case II, we presented the 5-fold cross-validation (CV) scores for the 200 replicates using boxplots in Figure 1 (a) and (b). The oracle (i.e., parameters are known) is compared with the RL-HMM, RL-HMM-fixed, and the RL model. In Figure 1 (a) for Case I, we observed that the RL-HMM demonstrates nearly the same performance as the oracle and outperforms RL-HMM-fixed, with RL showing the least favorable CV performance. In Figure 1 (b) for Case II, all models exhibit similar CV performance. On the other hand, recalling the estimated individual engagement probability for subject $i$ at trial $t$ as $\widehat{\gamma}_{i1t}=P(U_{it}=1\mid\mathbf{A}_{i},\widehat{\bm{\theta}})$, we estimate the decision-making strategy at trial $t$ for subject $i$ by $\widehat{U}_{it}=I(\widehat{\gamma}_{i1t}\geq 0.5)$. We present the averaged accuracy for estimating $U_{it}$ (i.e., $(nT)^{-1}\sum_{i=1}^{n}\sum_{t=1}^{T}I(\widehat{U}_{it}=U_{it})$) for the 200 replicates using boxplots in Figure 1 (c) and (d). In Figure 1 (c) for Case I, the RL-HMM outperforms RL-HMM-fixed, with RL showing the lowest accuracy because the RL model classifies every $U_{it}$ as ‘engaged’. In Figure 1 (d) for Case II, RL-HMM estimates more than $99.9\%$ of $U_{it}$ correctly for most of the replicates, demonstrating the robustness of RL-HMM when the model is misspecified. In the Supplementary Materials Section S.2, we employ additional simulations to demonstrate the robustness of our proposed method to the misspecification of the basis function $\bm{\phi}(a,s)$ and the choice of trend filtering order used to fit the nonparametric function $\zeta_{jt}$.
@@ 4.1 Probabilistic reward task and EMBARC study
In the EMBARC study (Trivedi et al.,, 2016), the PRT, as introduced by Pizzagalli et al., (2005), is a computer-based experiment used to assess subjects’ ability to adapt their behavior in response to rewards. During each trial of the task, participants are presented with a cartoon face displaying one of the two stimuli (i.e., a face with a short mouth or with a long mouth). The objective for participants is to identify the type of stimulus in each trial. Crucially, the PRT provides a higher frequency of rewarding participants for correctly identifying faces with a short mouth, designating it as the ‘rich reward’, while providing less frequent rewards for correctly identifying faces with a long mouth, referred to as the ‘lean reward’. Participants receive verbal instructions outlining the task’s goal, which is to maximize the total rewards. It is essential that participants understand that not all correct responses will be rewarded. To optimize their rewards, participants should focus on providing the correct response, regardless of the face associated with the higher reward. However, the difference in mouth length between the short and long mouths is intentionally kept small. The subtle difference in mouth length often leads to difficulties in accurately perceiving the stimuli presented, resulting in a tendency to prioritize states with higher rewards over those that are more accurate.
In the EMBARC study, each PRT session comprised 200 trials, divided into two blocks of 100 trials, with a 30-second break separating the blocks. In each block, 40 correct trials were followed by reward feedback. Correct identification of the rich stimulus was associated with significantly more positive feedback (30 out of 40 trials) compared to correct identification of the lean stimulus (10 out of 40 trials). The EMBARC study was designed to discern differences in reward learning abilities between patients diagnosed with Major Depressive Disorder (MDD) and a healthy control group. The PRT experiments involved 40 subjects from the control group and 168 subjects from the MDD group. Each subject in the MDD group had two PRT sessions, one session at the baseline before treatment (week 0) and one session after one week of treatment (week 1). Subjects in the control group also took two sessions in week 0 and week 1 (no treatment at both times). More details about the EMBARC study and the PRT experiment can be found in Guo et al., (2024).
@@ 4.2 Model fitting and results
We now apply the proposed methodology to our motivating PRT data. Let $\mathcal{S}=\{0,1\}$, where 00 and $1$ represent the type of stimulus (lean versus rich); $\mathcal{A}=\{0,1\}$, where 00 and $1$ represent the subject choosing lean and rich stimulus, respectively. We choose the basis $\phi(a,s)=(\delta_{00},\delta_{10},\delta_{01},\delta_{11})^{\top}$, where $\delta_{uv}=I(a=u,s=v)$ $u,v\in\{0,1\}$. We assume no initial preference for one stimulus state over the other at the beginning of each session. To ensure this, we specify the initial coefficients as $\bm{\alpha}=(\alpha/\rho)\cdot(1,0,0,1)$ and set the RL intercept values as $\nu_{0}=\nu_{1}=0$. In a preliminary analysis, we observed that the learning patterns may change between the first and second block. To mitigate any potential bias, we analyze the PRT data from the first block in this paper. We started by fitting a separate RL model for each session, using the method outlined in Section 2.1, and excluded sessions with a learning rate less than $10^{-3}$. We fitted our models for the MDD group that contains subjects diagnosed with MDD at week 0 (pre-treatment) and the control group with two repeated measurements, respectively. Based on the $5$-fold cross-validation likelihood score, we choose an RL-HMM with fixed transition probabilities over trials. This model has a higher score than an RL-HMM with time-varying transition probabilities or an RL-only model without decision-making strategy switching. The transition probabilities, $C_{01}$, are $2.8\%$ for the MDD group and $1.3\%$ for the control group. Similarly, the transition probabilities, $C_{11}$, are $99.2\%$ for the MDD group and $99.7\%$ for the control group. For inference, a nonparametric bootstrap was applied to generate 200 resampling sets for each group. Bootstrap standard errors and $95\%$ bootstrap confidence intervals are shown in Table 2.
@@
Table 2 reveals that the differences in the learning rate ($\beta$), reward sensitivity ($\rho$), and the initial probability of being in the ‘engaged’ strategy ($\pi_{1}$) between the MDD group and the control group are not statistically significant. This finding contrasts with previous research, such as Huys et al., (2013) and Guo et al., (2024), in which they reported a significantly lower reward sensitivity in the MDD group compared to the healthy control group in RL-only models. On the other hand, the MDD group presents a significantly lower initial coefficient ($\alpha$) compared to the control group. This indicates that the MDD group has a lower probability of selecting the correct state (i.e., distinguishing between long and short mouths) at the beginning of the task.
@@
Figure 2(a) presents the plug-in estimates of the individual engagement probability, $\widehat{\gamma}_{i1t}=P(U_{it}=1\mid\mathbf{A}_{i},\widehat{\bm{\theta}})$, for four randomly selected MDD patients which shows a general decreasing tendency of being engaged but with heterogeneous patterns between patients. The estimates and the $95\%$ pointwise bootstrap confidence intervals for the MDD and control groups engagement rates (defined in Section 2.2) are shown in Figure 2(b). We observe that the engagement rates for MDD and control groups are similar in the first 3/4 of the task. However, the MDD group engagement rate decreases and becomes smaller in the last 1/4 of the task compared to the control group engagement rate. This implies that individuals in the MDD group are more likely to be in the ‘lapse’ state after many trials in contrast to the control group. This finding suggests an alternative explanation for the reward processing abnormalities in MDD; it may be attributed to a reduced duration of making decisions using RL rather than a diminished sensitivity to rewards.
@@
To demonstrate the utilities of individual engagement scores (defined in Section 2.2), we consider its association with a commonly used clinical measure of depression severity, the Hamilton Depression Rating Scale (HDRS) (Hamilton,, 1960). In EMBARC, the HDRS was administered by trained clinicians to assess depression symptoms in different domains at the same time when the PRT was collected. One of items in HDRS assesses concentration (referred to as distraction level) on a liker scale of zero to three, where a higher value indicates greater difficulty in concentrating. In Figure 2(c), we present boxplots of individual engagement scores at different distraction levels (ranging from 0 to 3) in four periods of the task for patients in the MDD group. We observe that as the distraction level increases, the individual engagement score decreases. This effect is more pronounced in the first 3/4 of the trials. Furthermore, we performed a regression analysis of distraction levels on the individual engagement score using generalized linear models with binomial link functions. We found that the $p$-values for the coefficients are $0.004$, $0.043$, $0.013$, and $0.090$ for the four periods of the tasks, respectively. This implies that a lack of engagement, particularly at the beginning of the task, is associated with difficulties in concentration.
@@
Define $\widehat{U}_{it}=I(\widehat{\gamma}_{i1t}\geq 0.5)$ to be the predicted type of decision-making strategy for the subject $i$ at trial $t$. The correct response rates (i.e. the accurate identification of short and long mouth stimuli) are $80.9\%$ and $48.4\%$ for the ‘engaged’ and ‘lapse’ state within the MDD group, and $84.3\%$ and $55.3\%$ for the control group. This suggests that the MDD group has a lower correct response rate in the ‘engaged’ state, and the correct response rate in the ‘lapse’ state is close to random guessing in both groups. Additionally, Iigaya et al., (2018) showed that different decision-making strategies may result in different lengths of intertrial intervals (ITIs), which is the response time between the presence of a stimulus and the decision making in a trial. Furthermore, we obtain the smoothed ITIs over trials for the two decision-making strategies for two groups separately using local polynomial regression (loess function in R). The mean estimates and their $95\%$ confidence intervals for the mean curves are displayed in Figure 2(d). It reveals that, on average, the ‘engaged’ strategy requires more time for individuals to make decisions compared to the ‘lapse’ strategy. Meanwhile, the control group, on average, takes less time to make decisions than the MDD group.